OpenStudy (anonymous):

(2 cos x+3)(2 sin x-1)=0

OpenStudy (anonymous):

2sinx - 1 = 0 2sinx = 1 sin x = 1/2 x = pi/6

OpenStudy (anonymous):

On the other side for 2cosx + 3 = 0 cosx = -3/2 \[x = 2k \pi \pm \arccos(-3/2)\]

OpenStudy (anonymous):

Oh, and in the end you can combine the two results so that \[x \in \left\{ \frac{ \pi }{ 6 }\ , 2k \pi \pm acrccos(-\frac{ 3 }{ 2 }), k \in \mathbb{Z} \right\}\]